OpenStudy (anonymous):

3 • 5 = 5 • 3 is an example of which algebraic property?

OpenStudy (zzr0ck3r):

commutative property of multiplication a*b = b*a

OpenStudy (anonymous):

5 + a = b is the same as b = 5 + a. This is an example of which algebraic property?